# Knowledge Mechanisms in Large Language Models: A Survey and Perspective

### v1

@@ 4.2 Group Evolution
Besides individual learning, social interaction plays a pivotal role in the acquisition of new knowledge and is a key driver of human societal development Baucal et al. (2014). LLMs, also known as agents, collaborate to accomplish complex tasks during group evolution, each bearing unique knowledge that may sometimes contradict each other. Therefore, contrary to individual evolution, group evolution encounters intensified conflicts, such as conflicts in specialized expertise among agents, competing interests, cultural disparities, moral dilemmas, and others. To achieve consensus and resolve conflicts, agents must first clarify their own and others’ goals (beliefs) through internal representations in models Zhu et al. (2024); Zou et al. (2023). Agents then discuss, debate, and reflect on shared knowledge through various communication methods Soltoggio et al. (2024), e.g., prompt instructions, task and agent descriptions, parameter signals (activation and gradient), and representations of models. However, conformity of agents, which tends to believe the majority’s incorrect answers rather than maintaining their own, hinders conflict resolution during group evolution Zhang et al. (2023a).
Note that the group also struggles with automating moral decision-making when facing moral conflicts. Specifically, agents in the group miss ground truth for moral “correctness” and encounter dilemmas due to changes in moral norms over time Hagendorff and Danks (2023). Generally, when, what, and how to share knowledge in the communication process to maximize learning efficiency and long-term expectations are still open questions in group evolution.
@@ 4.3 Comparison of Different Evolution Strategies
Individuals and groups achieve dynamic intelligence primarily through two strategies: updating internal parametric knowledge Zhou et al. (2023a); Qiao et al. (2024) and leveraging external knowledge 444 Leveraging external knowledge includes using prompts Xie et al. (2024), ICL, and RAG. Huang and Huang (2024); Xie et al. (2024). These two strategies are usually used together in applications Yang et al. (2024b).
@@
Updating internal parametric knowledge necessitates high-quality data for parameter adjustments Vashishtha et al. (2024). Data proves pivotal when fine-tuning models to acquire new knowledge. Ovadia et al. (2024) also posit that the continued training of LLMs via unsupervised tuning generally exhibits suboptimal performance when it comes to acquiring new knowledge. Note that updating internal parametric knowledge requires resolving conflicts among internal parameters. The crux of effective internal knowledge updating lies in preserving the consistency of the model’s parameter knowledge before and after tuning. In contrast, leveraging external knowledge requires managing conflicts within the external knowledge itself  [^5] Inconsistencies in external information are common, as external documents often contain conflicting data, particularly in contexts for RAG. as well as conflicts between external and internal knowledge Xu et al. (2024b); Liu et al. (2024a). Besides, parametric knowledge compresses extensive information, promoting grokking and enhancing generalization Wang et al. (2024a). In contrast, leveraging external knowledge avoids high training costs but necessitates substantial maintenance and retrieval costs for every user query. Therefore, the combination of these two strategies is promising. An attempt for combination Yang et al. (2024b) suggests employing RAG for low-frequency knowledge and parametric strategy for high-frequency knowledge.
@@ 5.1 What Knowledge Have LLMs Learned?
Critics question whether LLMs truly have knowledge or if they are merely mimicking Schwarzschild et al. (2024), akin to the “Stochastic Parro” Bender et al. (2021) and “Clever Hans” Shapira et al. (2024). We first review the doubts from the following three levels through observation phenomena: 1) Memorization: LLMs primarily rely on positional information over semantic understanding Li et al. (2022) to predict answers. Additionally, LLMs may generate different answers for the same question due to different expressions. 2) Comprehension and application: Allen-Zhu and Li (2023b) argue that LLMs hardly efficiently apply knowledge from pre-training data, even when such knowledge is perfectly stored and fully extracted from LLMs. Therefore, LLMs struggle with diverse reasoning tasks Nezhurina et al. (2024); Gutiérrez et al. (2024) as well as the reverse curse Berglund et al. (2023). Besides, LLMs are not yet able to reliably act as text world simulators and encounter difficulties with planning Wang et al. (2024d). 3) Creation: Although LLMs are capable of generating new terms, their quality often falls below that created by humans Raiola (2023). Even though LLMs possess knowledge, some critics argue that current analysis methods may only explain low-level co-occurrence patterns, not internal mechanisms. The primary criticism asserts that the components responsible for certain types of knowledge in LLM fail to perform effectively in practical applications Hase et al. (2023). In addition, the components responsible for specific knowledge within LLMs vary under different methods. For these criticisms, Chen et al. (2024f, d) propose degenerate neurons and posit that different degenerate components indeed independently express a fact. Chen et al. (2024e) delineate the differences in the mechanisms of knowledge storage and representation, proposing the Query Localization Assumption to response these controversies. Zhu and Li (2023) further observe that knowledge may be memorized but not extracted due to the knowledge not being sufficiently augmented (e.g., through paraphrasing, sentence shuffling) during pretraining. Hence, rewriting the training data to provide knowledge augmentation and incorporating more instruction fine-tuning data in the pre-training stage can effectively alleviate the above challenges and criticisms.
@@
Despite considerable criticism, the mainstream view is that current LLMs may possess basic world knowledge but hardly master principle knowledge of underlying rules for reasoning and creativity. In other words, LLMs master knowledge memorization (discussed in §3.1). Although LLMs possess the foundational ability to comprehend and apply knowledge (discussed in §3.2), they still struggle with reasoning and planning in complex tasks due to the fragility of knowledge in LLMs (elaborated in §5.2). These reasoning and planning abilities usually require to be induced through techniques such as ICL and CoT. Unfortunately, current LLMs are nearly incapable of creation due to the architectural limitations (discussed in §3.3). Therefore, some scholars explore various architectural choices (e.g., Mamba) and training procedures. To build a transparent and trustworthy model Yu et al. (2023c), several initiatives suggest encouraging sparsity, modularity, and monosemanticity in network architectures. Besides, recent research attempts to manipulate neurons, knowledge circuits, or representations Allen-Zhu and Li (2023b); Zou et al. (2023); Wu et al. (2024); Li et al. (2023a) to explore more knowledge and awaken the reasoning and planning capabilities of LLMs.
Remarks: LLMs possess basic world knowledge but encounter challenges in application (§5.2). Some knowledge still lies beyond their current cognitive scope (§5.3). For example, due to architectural limitations, LLMs struggle with creativity.
5.2 Improper Learning or Misuse?
@@
The knowledge learned by LLMs is fragile, leading to issues of hallucination, knowledge conflicts, and secure risk  [^6] The secure risk is elaborated in §6.4.. Hallucination denotes content generated by LLMs that diverges from real-world facts or inputs Huang et al. (2023b); Xu et al. (2024d); Farquhar et al. (2024); Chen et al. (2024c). On the one hand, factuality hallucination underscores the disparity between generated content and real-world knowledge. On the other hand, faithfulness hallucination describes the departure of generated content from user instructions or input context, as well as the coherence maintained within the generated content. Knowledge Conflict inherently denotes inconsistencies in knowledge Xu et al. (2024b); Kortukov et al. (2024). On the one hand, internal memory conflicts within the model cause LLMs to exhibit unpredictable behaviors and generate differing results to inputs which are semantically equivalent but syntactically distinct Xu et al. (2024b); Wang et al. (2023a); Feng et al. (2023b); Raj et al. (2022). On the other hand, context-memory conflict emerges when external context knowledge contradicts internal parametric knowledge Jin et al. (2024b); Yao et al. (2024); Hoffmann et al. (2022); Kasai et al. (2023).
@@
We posit that these challenges mainly derive from improper learning data. Specifically, hallucination is introduced by data Kang and Choi (2023); Weng (2024), heightened during the pre-training Brown et al. (2020); Chiang and Cholak (2022), alignment Azaria and Mitchell (2023); Ouyang et al. (2022), and deficiencies in decoding strategies Fan et al. (2018); Chuang et al. (2023); Shi et al. (2023). Internal memory conflict can be attributed to training corpus bias Wang et al. (2023b), and exacerbated by decoding strategies Lee et al. (2022b) and knowledge editing. Context-memory conflict arises mainly from the absence of accurate knowledge during training, necessitating retrieval from databases and the Web. Wang et al. (2024a) further delve into the mechanism between parametric knowledge and learning data, demonstrate that training data distribution qualitatively influences generalization behavior Jiang et al. (2024a) and how LLMs encapsulate or enhance knowledge related to reasoning into specific components, e.g., bridge layer Wang et al. (2024a). Similar mechanism analysis also supports the above conclusion, indicating that hallucinations arise from a lack of mover heads Yao et al. (2024); Yu et al. (2024b), while knowledge conflicts stem from circuit competition failure in the last few layers Lv et al. (2024); Merullo et al. (2023b); Hase et al. (2023); Ju et al. (2024); Jin et al. (2024c). Additionally, data quantity is crucial for knowledge robustness. Specifically, LLMs can systematically learn comprehensive understandings of the world from extensive datasets, while little data during post-training stage may compromise the robustness of knowledge representation. This assumption is confirmed by numerous failures of post-training. For example, SFT exacerbates hallucinations Gekhman et al. (2024); Kang et al. (2024), and knowledge editing amplifies knowledge conflicts Li et al. (2023c); Yang et al. (2024c). Note that safety issues usually caused by the distribution of unseen data (adversarial input) Wei et al. (2023); Li et al. (2024b), which is elaborated in §6.4.
@@
5.3 Dark Knowledge Hypothesis for LLMs?
@@
We hypothesize that there will still exist dark knowledge for intelligence in the future. As shown in Fig 4, dark knowledge describes knowledge unknown to human or machine from the following three situations: 1) knowledge unknown to human & known to machine (UH, KM). Machines leverage vast amounts of data to explore internal patterns, whereas humans struggle with processing such data due to physiological limitations on data processing capacity and computational limits Burns et al. (2023); McAleese et al. (2024). (UH, KM) includes gene prediction, intelligent transportation systems, and more. Specifically, the structural elucidation of proteins remains mysterious to humans for a long time. Cryo-electron microscopy, through capturing millions of images, first reveals the three-dimensional structures of proteins. Now, neural models can directly predict protein properties with high efficiency and accuracy Pak et al. (2023). 2) knowledge known to human & unknown to machine (UH, KM). On the one hand, some scholars claim that machine can possess a “Theory of Mind” capability Zhu et al. (2024) and emotions Normoyle et al. (2024). On the other hand, critics contend that machine lacks sentient and merely probabilistically generates tokens Wang et al. (2024d). Specifically, some abstract knowledge like hunger, happiness, loneliness, and others cannot be fully represented by discrete data points. The above knowledge akin to NP-hard problems Hochba (1997), is also exceedingly complex and nearly impossible to model perfectly with current algorithms. Therefore, opponents argue that no matter how many parameters machine possesses, it cannot learn all the knowledge that human has mastered. 3) knowledge unknown to human & unknown to machine (UH, UM) is beyond our cognition. Generally, Dark knowledge extends beyond current data and model architectures Tseng et al. (2024). (UH, UM) necessitates human-machine collaboration. Yet, there is no definitive conclusion on whether (UH, KM) and (KH, UM) will be solved by model architecture, training data, and computational resources. Note that plain knowledge known to human and machine in Fig 4 encompasses well-defined historical events, mathematical theorems, physical laws, etc.
@@
5.4 How to Benefit from Interdisciplinarity?
How can LLMs continuously narrow the boundaries of dark knowledge and achieve higher level intelligence by leveraging the human experience of perpetual knowledge exploration throughout history? We draw inspirations from the following interdisciplinary studies.
Neuroscience studies the structure and function of the brain at molecular, cellular, neural circuit, and neural network levels Squire et al. (2012). Generally, both mechanism analysis in LLMs and neuroscience utilize observation and intervention methods to investigate the basic principles of knowledge learning and memory, decision-making, language, perception, and consciousness. The biological signals of the human brain and the internal activation signals in LLMs are capable of reciprocal transformation Caucheteux et al. (2023); Feng et al. (2023a); Mossing et al. (2024); Flesher et al. (2021). Benefiting from advancements in neuroscience Jamali et al. (2024); de Schotten et al. (2022); Lee et al. (2022a), mechanism analysis in LLMs has identified analogous function neurons and regions Zhao et al. (2023a), and knowledge circuits Yao et al. (2024). Furthermore, leveraging plasticity theory in neuroscience, LLMs explain the underlying technical support for intelligence Sæbø and Brovold (2024). In the future, mechanism analysis of LLMs may draw inspirations from neuroscience, guiding the next generation of artificial intelligence in organizing neural frameworks and in the storage and utilization of knowledge Momeni et al. (2024); Yang et al. (2024b).
Cognitive Science focuses on the mind and its processes Kolak et al. (2006); Baronchelli et al. (2013), which include language, perception, memory, attention, reasoning, emotion and mental state. Although cognitive science and neuroscience overlap in their research content, cognitive science focuses more on abstract knowledge such as mental states and emotions rather than specific knowledge. Therefore, Zhu et al. (2024) track beliefs of self and others (formulated as “Theory of Mind”) in LLMs from the psychological perspective within cognitive science. Wang et al. (2022) further observe social-cognitive skill in multi-agent communication and cooperation. Generally, there is potential to explore advanced cognitive capabilities in LLMs from the perspective of cognitive science.

### v2

@@ 4.2 Group Evolution
Besides individual learning, social interaction plays a pivotal role in the acquisition of new knowledge and is a key driver of human societal development Baucal et al. (2014). LLMs, also known as agents, collaborate to accomplish complex tasks during group evolution, each bearing unique knowledge that may sometimes contradict each other. Therefore, contrary to individual evolution, group evolution encounters intensified conflicts, such as conflicts in specialized expertise among agents, competing interests, cultural disparities, moral dilemmas, and others. To achieve consensus and resolve conflicts, agents must first clarify their own and others’ goals (beliefs) through internal representations in models Zhu et al. (2024); Zou et al. (2023). Agents then discuss, debate, and reflect on shared knowledge through various communication methods Soltoggio et al. (2024), e.g., prompt instructions, task and agent descriptions, parameter signals (activation and gradient), and representations of models. However, conformity of agents, which tends to believe the majority’s incorrect answers rather than maintaining their own, hinders conflict resolution during group evolution Zhang et al. (2023a). Note that the group also struggles with automating moral decision-making when facing moral conflicts. Specifically, agents in the group miss ground truth for moral “correctness” and encounter dilemmas due to changes in moral norms over time Hagendorff and Danks (2023). Generally, when, what, and how to share knowledge in the communication process to maximize learning efficiency and long-term expectations are still open questions in group evolution.
@@ 4.3 Comparison of Different Evolution Strategies
Individuals and groups achieve dynamic intelligence primarily through two strategies: updating internal parametric knowledge Zhou et al. (2023a); Qiao et al. (2024) and leveraging external knowledge 555 Leveraging external knowledge includes using prompts Xie et al. (2024), ICL, and RAG. Huang and Huang (2024); Xie et al. (2024). These two strategies are usually used together in applications Yang et al. (2024b).
@@
Updating internal parametric knowledge necessitates high-quality data for parameter adjustments Vashishtha et al. (2024); Cao et al. (2024a). Data proves pivotal when fine-tuning models to acquire new knowledge. Ovadia et al. (2024) also posit that the continued training of LLMs via unsupervised tuning generally exhibits suboptimal performance when it comes to acquiring new knowledge. Note that updating internal parametric knowledge requires resolving conflicts among internal parameters. The crux of effective internal knowledge updating lies in preserving the consistency of the model’s parameter knowledge before and after tuning. In contrast, leveraging external knowledge requires managing conflicts within the external knowledge itself  [^6] Inconsistencies in external information are common, as external documents often contain conflicting data, particularly in contexts for RAG. as well as conflicts between external and internal knowledge Xu et al. (2024b); Liu et al. (2024a). Besides, parametric knowledge compresses extensive information, promoting grokking and enhancing generalization Wang et al. (2024a). In contrast, leveraging external knowledge avoids high training costs but necessitates substantial maintenance and retrieval costs for every user query. Therefore, the combination of these two strategies is promising. An attempt for combination Yang et al. (2024b) suggests employing RAG for low-frequency knowledge and parametric strategy for high-frequency knowledge.
5 Application of Knowledge Mechanism
The mechanism analysis of knowledge utilization and evolution may provide an avenue to construct more efficient and trustworthy models in practice.
5.1 Efficient LLMs
Researchers have been working to reduce the cost of training and inference for LLMs through various optimization strategies, including architecture Ainslie et al. (2023); Fedus et al. (2022), data quality Kaddour (2023), parallelization Qi et al. (2024), generalization theory Zhang et al. (2024d), hardware Dey et al. (2023), scaling laws Hoffmann et al. (2022), optimizer Liu et al. (2023a), etc. The underlying knowledge mechanisms offer LLMs new potential for efficiently storing, utilizing, and evolving knowledge.
For knowledge storage and utilization in LLMs, knowledge (memory) circuit provides the theory to decompose the knowledge computations of an LLM into smaller, recurring parts Yang et al. (2024b). These smaller parts guide the determination of which types of knowledge should be encoded into parameters. Therefore, $\text{Memory}^{3}$ Yang et al. (2024b) designs an explicit memory mechanism for Transformer-based LLMs, alleviating the burden of parameter size. Specifically, $\text{Memory}^{3}$ designs external information, explicit memory, and implicit memory for different usage frequencies, reducing writing and reading costs. For knowledge evolution, the knowledge mechanism analysis inspires editing and model merging. The details of editing technologies can be found in §4.2. Model merging technologies  [^7] Model merging also includes methods that directly interpolation Goddard et al. (2024) or randomly fusing Yu et al. (2023b), ignoring parameter directions. These methods are naive and are not the focus of our discussion here. leverage parameter directions to combine multiple task-specific models into a single multitask model without performing additional training rather than training from scratch. For instance, Task Arithmetic Ilharco et al. (2023) identifies the weight directions of task capabilities in different models, and then integrates a more powerful model by arithmetic operations on weight directions. TIES Yadav et al. (2023) resolves parameters directions conflicts, and merges only the parameters that are in alignment with the final agreed-upon sign. Akiba et al. (2024) further propose evolutionary optimization of model merging, which automatically discovers effective combinations of open-source models, harnessing their group intelligence without requiring extensive training data or computational resources. Besides, the Lottery Ticket Hypothesis Frankle and Carbin (2019) provides a cornerstone for model compression, generalizing across various datasets, optimizers, and model architectures Morcos et al. (2019); Chen et al. (2021). However, model compression often limits the success of editing and model merging Kolbeinsson et al. (2024). This phenomenon poses challenges for practical implementations, highlighting the need for more effective strategies.
5.2 Trustworthy LLMs
Numerous studies investigate the underlying causes of security risks. In particular, Wei et al. (2023) delve into the safety of LLM and reveal that the success of jailbreak is mainly due to the distribution discrepancies between malicious attacks and training data. Geva et al. (2022b) and Wang et al. (2024b) further discover that some parameters within LLMs, called toxic regions, are intrinsically tied to the generation of toxic content. Ji et al. (2024a) even conjecture that LLMs resist alignment. Therefore, traditional aligned methods, DPO Rafailov et al. (2023) and SFT, seem to merely bypass toxic regions Lee et al. (2024); Wang et al. (2024b), making them susceptible to other jailbreak attacks Zhang et al. (2023d).
Inspired by the knowledge mechanism analysis in LLMs, a promising trustworthy strategy may be designing architecture and training process during the pre-training phase to encouraging modularity Liu et al. (2024c, 2023b), sparsity Chughtai et al. (2023), and monosemanticity Bricken et al. (2023); Jermyn et al. (2022), which make the reverse engineering process more tractable Jermyn et al. (2022); Bricken et al. (2023); Liu et al. (2024c); Tamkin et al. (2023). Yet, maintaining sparsity for a vast amount of world knowledge requires substantial resources, and whether monosemantic architecture can support advanced intelligence remains elusive. Besides, machine unlearning Nguyen et al. (2022); Tian et al. (2024) aims to forget privacy or toxic information learned by LLMs. However, these unlearning methods suffer overfitting, forgetting something valuable due to the difficulty of disentangling verbatim memorization and general capabilities Huang et al. (2024c). Another alternative technique is knowledge editing, precisely modifying LLMs using few instances during the post-training stage Mazzia et al. (2023); Yao et al. (2023); Wang et al. (2023d); Hase et al. (2024); Qian et al. (2024b). Extensive experiments demonstrate that knowledge editing has the potential to detoxify LLMs Yan et al. (2024). Specifically, Wu et al. (2023) and Geva et al. (2022b) deactivate the neurons related to privacy information and toxic tokens, respectively. Wang et al. (2024b) identify and then erases toxic regions in LLMs. However, knowledge editing also introduces side effects, such as the inability of the modified knowledge to generalize to multi-hop tasks Zhong et al. (2023); Li et al. (2023d); Cohen et al. (2023); Kong et al. (2024) and the potential to impair the model’s general capabilities Gu et al. (2024); Qin et al. (2024). Therefore, recent efforts focus on representation editing instead of editing parameters in knowledge editing Zou et al. (2023); Turner et al. (2023); Zhou et al. (2023b); Zhu et al. (2024). These representations (hidden states) within LLMs can track and address a wide range of safety-relevant problems, including honesty, harmlessness, and power seeking. Later, Wu et al. (2024) develop a family of representation finetuning methods to update new knowledge. Zou et al. (2024) propose circuit-breaking Li et al. (2023b), directly controlling the representations that are responsible for harmful outputs. However, these representation editing strategies require meticulous hyperparameter tuning for each task. More efficient optimization methods are needed to align with computational or temporal constraints.
@@ 6 Discussion
In this section, we discuss some open questions and seek to explore their essence and underlying principles. Specifically, we discuss what knowledge LLMs have learned in §6.1, examine the fragility of the learned knowledge in application in §6.2, analyze the dark knowledge not yet learned by machines or humans in §6.3, and explore how LLMs can expand the boundaries of unknown knowledge from interdisciplinary perspectives §6.4.
@@ 6.1 What Knowledge Have LLMs Learned?
Critics question whether LLMs truly have knowledge or if they are merely mimicking Schwarzschild et al. (2024), akin to the “Stochastic Parro” Bender et al. (2021) and “Clever Hans” Shapira et al. (2024). We first review the doubts from the following three levels through observation phenomena: 1) Memorization: LLMs primarily rely on positional information over semantic understanding Li et al. (2022) to predict answers. Additionally, LLMs may generate different answers for the same question due to different expressions. 2) Comprehension and application: Allen-Zhu and Li (2023b) argue that LLMs hardly efficiently apply knowledge from pre-training data, even when such knowledge is perfectly stored and fully extracted from LLMs. Therefore, LLMs struggle with diverse reasoning tasks Nezhurina et al. (2024); Gutiérrez et al. (2024) as well as the reversal curse Berglund et al. (2023). Besides, LLMs are not yet able to reliably act as text world simulators and encounter difficulties with planning Wang et al. (2024d). 3) Creation: Although LLMs are capable of generating new terms, their quality often falls below that created by humans Raiola (2023). Even though LLMs possess knowledge, some critics argue that current analysis methods may only explain low-level co-occurrence patterns, not internal mechanisms. The primary criticism asserts that the components responsible for certain types of knowledge in LLM fail to perform effectively in practical applications Hase et al. (2023). In addition, the components responsible for specific knowledge within LLMs vary under different methods. For these criticisms, Chen et al. (2024f, d) propose degenerate neurons and posit that different degenerate components indeed independently express a fact. Chen et al. (2024e) delineate the differences in the mechanisms of knowledge storage and representation, proposing the Query Localization Assumption to response these controversies. Zhu and Li (2023) further observe that knowledge may be memorized but not extracted due to the knowledge not being sufficiently augmented (e.g., through paraphrasing, sentence shuffling) during pretraining. Hence, rewriting the training data to provide knowledge augmentation and incorporating more instruction fine-tuning data in the pre-training stage can effectively alleviate the above challenges and criticisms.
@@
Despite considerable criticism, the mainstream view is that current LLMs may possess basic world knowledge but hardly master principle knowledge of underlying rules for reasoning and creativity. In other words, LLMs master knowledge memorization (discussed in §3.1). Although LLMs possess the foundational ability to comprehend and apply knowledge (discussed in §3.2), they still struggle with reasoning and planning in complex tasks due to the fragility of knowledge in LLMs (elaborated in §6.2). These reasoning and planning abilities usually require to be induced through techniques such as ICL and CoT. Unfortunately, current LLMs are nearly incapable of creation due to the architectural limitations (discussed in §3.3). Therefore, some scholars explore various architectural choices (e.g., Mamba Gu and Dao (2023)) and training procedures. Besides, recent research attempts to manipulate neurons, knowledge circuits, or representations Allen-Zhu and Li (2023b); Zou et al. (2023); Wu et al. (2024); Li et al. (2023a) to explore more knowledge and awaken the reasoning and planning capabilities of LLMs.
Remarks: LLMs have learned basic knowledge of the world by momorization. However, the learned knowledge is fragile, leading to challenges in knowledge comprehension and application. Unfortunately, due to architectural limitations, current LLMs struggle with creation.
6.2 Why Is Learned Knowledge Fragile?
@@
The knowledge learned by LLMs is fragile, leading to challenges in application including hallucination, knowledge conflicts, and secure risk  [^8] The secure risk is elaborated in §5.2.. Hallucination denotes content generated by LLMs that diverges from real-world facts or inputs Huang et al. (2023b); Xu et al. (2024d); Farquhar et al. (2024); Chen et al. (2024c). On the one hand, factuality hallucination underscores the disparity between generated content and real-world knowledge. On the other hand, faithfulness hallucination describes the departure of generated content from user instructions or input context, as well as the coherence maintained within the generated content. Knowledge Conflict inherently denotes inconsistencies in knowledge Xu et al. (2024b); Kortukov et al. (2024). On the one hand, internal memory conflicts within the model cause LLMs to exhibit unpredictable behaviors and generate differing results to inputs which are semantically equivalent but syntactically distinct Xu et al. (2024b); Wang et al. (2023a); Feng et al. (2023b); Raj et al. (2022). On the other hand, context-memory conflict emerges when external context knowledge contradicts internal parametric knowledge Xu et al. (2024b); Mallen et al. (2023).
@@
We posit that these challenges mainly derive from improper learning data. Specifically, hallucination is introduced by data Kang and Choi (2023); Weng (2024), heightened during the pre-training Brown et al. (2020); Chiang and Cholak (2022), alignment Azaria and Mitchell (2023); Ouyang et al. (2022), and deficiencies in decoding strategies Fan et al. (2018); Chuang et al. (2023); Shi et al. (2023). Internal memory conflict can be attributed to training corpus bias Wang et al. (2023b), and exacerbated by decoding strategies Lee et al. (2022b) and knowledge editing. Context-memory conflict arises mainly from the absence of accurate knowledge during training, necessitating retrieval from databases and the Web. Wang et al. (2024a) and Antoniades et al. (2024) further delve into the mechanism between parametric knowledge and learning data, demonstrate that training data distribution qualitatively influences generalization behavior Jiang et al. (2024a) and how LLMs encapsulate or enhance knowledge related to reasoning into specific components, e.g., bridge layer Wang et al. (2024a). Similar mechanism analysis also supports the above conclusion, indicating that hallucinations arise from a lack of mover heads Yao et al. (2024); Yu et al. (2024b), while knowledge conflicts stem from circuit competition failure in the last few layers Lv et al. (2024); Merullo et al. (2023b); Hase et al. (2023); Ju et al. (2024); Jin et al. (2024b). Additionally, data quantity is crucial for knowledge robustness. Specifically, LLMs can systematically learn comprehensive understandings of the world from extensive datasets, while little data during post-training stage may compromise the robustness of knowledge representation. This assumption is confirmed by numerous failures of post-training. For example, SFT exacerbates hallucinations Gekhman et al. (2024); Kang et al. (2024), and knowledge editing amplifies knowledge conflicts Li et al. (2023d); Yang et al. (2024c). Note that safety issues usually caused by the distribution of unseen data (adversarial input) Wei et al. (2023); Li et al. (2024b), which is elaborated in §5.2.
@@
6.3 Does Difficult-to-Learn “Dark Knowledge” Exist?
@@
We hypothesize that there will still exist dark knowledge for intelligence in the future. As shown in Fig 4, dark knowledge describes knowledge unknown to human or machine from the following three situations: 1) knowledge unknown to human & known to machine (UH, KM). Machines leverage vast amounts of data to explore internal patterns, whereas humans struggle with processing such data due to physiological limitations on data processing capacity and computational limits Burns et al. (2023); McAleese et al. (2024). (UH, KM) includes gene prediction, intelligent transportation systems, and more. Specifically, the structural elucidation of proteins remains mysterious to humans for a long time. Cryo-electron microscopy, through capturing millions of images, first reveals the three-dimensional structures of proteins. Now, neural models can directly predict protein properties with high efficiency and accuracy Pak et al. (2023). 2) knowledge known to human & unknown to machine (UH, KM). On the one hand, some scholars claim that machine can possess a “Theory of Mind” capability Zhu et al. (2024) and emotions Normoyle et al. (2024). On the other hand, critics contend that machine lacks sentience Alvero and Peña (2023) and merely probabilistically generates tokens. The causes, extent, and dynamics of these emotions and sentience (like hunger, happiness, and loneliness) are subtle and intricate, making precise mathematical modeling by the machine exceptionally challenging. Specifically, different factors are tightly coupled, making it nearly impossible to disentangle clear input-output relationships as with well-defined factual knowledge. The sentient knowledge also exhibits chaotic behavior Li et al. (2020); Debbouche et al. (2021), being highly sensitive to initial conditions, where small changes can lead to vastly different outcomes Segretain et al. (2020). Therefore, opponents argue that no matter how many parameters machine possesses, it cannot learn all the knowledge that human has mastered. 3) knowledge unknown to human & unknown to machine (UH, UM) is beyond our cognition, e.g., the uncertainty in quantum mechanics and the origin of the universe. Generally, Dark knowledge extends beyond current data and model architectures Tseng et al. (2024). (UH, UM) necessitates human-machine collaboration. Yet, there is no definitive conclusion on whether (UH, KM) and (KH, UM) will be solved by model architecture, training data, and computational resources. Note that plain knowledge known to human and machine in Fig 4 encompasses well-defined historical events, mathematical theorems, physical laws, etc.
@@
6.4 How to Explore More Knowledge from Interdisciplinary Inspiration?
How can LLMs continuously narrow the boundaries of dark knowledge and achieve higher level intelligence by leveraging the human experience of perpetual knowledge exploration throughout history? We may draw inspirations from the following interdisciplinary studies.
Neuroscience studies the structure and function of the brain at molecular, cellular, neural circuit, and neural network levels Squire et al. (2012). Generally, both mechanism analysis in LLMs and neuroscience utilize observation and intervention methods to investigate the basic principles of knowledge learning and memory, decision-making, language, perception, and consciousness. The biological signals of the human brain and the internal activation signals in LLMs are capable of reciprocal transformation Caucheteux et al. (2023); Feng et al. (2023a); Mossing et al. (2024); Flesher et al. (2021). Benefiting from advancements in neuroscience Jamali et al. (2024); de Schotten et al. (2022); Lee et al. (2022a), mechanism analysis in LLMs has identified analogous function neurons and regions Zhao et al. (2023a), and knowledge circuits Yao et al. (2024). Besides, leveraging plasticity theory in neuroscience, LLMs explain the underlying technical support for intelligence Sæbø and Brovold (2024). In the future, mechanism analysis of LLMs may draw inspirations from neuroscience, guiding the next generation of artificial intelligence in organizing neural frameworks and in the storage and utilization of knowledge Momeni et al. (2024); Yang et al. (2024b).
Cognitive Science focuses on the mind and its processes Kolak et al. (2006); Baronchelli et al. (2013), which include language, perception, memory, attention, reasoning, emotion and mental state. Although cognitive science and neuroscience overlap in their research content, cognitive science focuses more on abstract knowledge such as mental states and emotions rather than specific knowledge. Therefore, Zhu et al. (2024) track beliefs of self and others (formulated as “Theory of Mind”) in LLMs from the psychological perspective within cognitive science. Wang et al. (2022) further observe social-cognitive skill in multi-agent communication and cooperation. Generally, there is potential to explore advanced cognitive capabilities in LLMs from the perspective of cognitive science Vilas et al. (2024).